OpenStudy (anonymous):

Find the product of (x + 7)2. x2 − 14x + 49 x2 + 14x + 49 x2 − 49 x2 + 49

Parth (parthkohli):

\[ \large \begin{aligned} (x + 7)^2&=(x + 7)(x+7) \\ &=x(x + 7) + 7(x + 7)\end{aligned}\]Now, use your knowledge of the distributive property.

OpenStudy (anonymous):

OH OKAY ILL USE MY KNOWLEDGE

OpenStudy (anonymous):

thanks